# Toward Relative Positional Encoding in Spiking Transformers

### v2

@@ Abstract
Spiking neural networks (SNNs) are bio-inspired networks that mimic how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities. SNNs with self-attention mechanisms (spiking Transformers) have recently shown great advancements in various tasks, and inspired by traditional Transformers, several studies have demonstrated that spiking absolute positional encoding can help capture sequential relationships for input data, enhancing the capabilities of spiking Transformers for tasks such as sequential modeling and image classification. However, how to incorporate relative positional information into SNNs remains a challenge. In this paper, we introduce several strategies to approximate relative positional encoding (RPE) in spiking Transformers while preserving the binary nature of spikes. Firstly, we formally prove that encoding relative distances with Gray Code ensures that the binary representations of positional indices maintain a constant Hamming distance whenever their decimal values differ by a power of two, and we propose Gray-PE based on this property. In addition, we propose another RPE method called Log-PE, which combines the logarithmic form of the relative distance matrix directly into the spiking attention map. Furthermore, we extend our RPE methods to a two-dimensional form, making them suitable for processing image patches. We evaluate our RPE methods on various tasks, including time series forecasting, text classification, and patch-based image classification, and the experimental results demonstrate a satisfying performance gain by incorporating our RPE methods across many architectures. Our results provide fresh perspectives on designing spiking Transformers to advance their sequential modeling capability, thereby expanding their applicability across various domains.
@@ 3.2 Spiking Self-Attention
As a result, the attention map $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers. The outputs of the SSA, as well as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, are all spike matrices containing only values of 00 and $1$. The parameters $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}$ are all learnable parameters.
@@ 3.3 Relative Positional Encoding
Here, $\mathbf{R}_{i,j}$ represents the relative positional bias between the $i$-th query and the $j$-th key positions. A critical aspect of RPE is its adherence to distance consistency: the magnitude of $\mathbf{R}_{i,j}$ is determined exclusively by the relative positional offset $|i-j|$, ensuring that the model systematically differentiates between proximal and distant tokens. This property enhances the model’s capacity to capture long-range dependencies and generalize across variations in sequence length and structure.
@@ 3.4 Hamming Distance
$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),\quad\text{where}\quad\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (6)
@@ 3.5 Gray Code
$\small G(x)=x\oplus\left(x\gg 1\right),$ (7)
@@ 4.3 Gray-PE
We propose that the Gray Code can serve as an approximate approach to relative positional encoding for spiking Transformers. This is supported by the following Theorem 1:
@@
As illustrated in Figure 2 (b), the Hamming distance $d_{H}(0,1)$ and $d_{H}(1,2)$ both equal $1$ because their relative distance is $1$, i.e., $02^{n},n=$ 0. Similarly, $d_{H}(0,2)=d_{H}(1,3)$, and $d_{H}(0,4)=d_{H}(1,5)$, as their relative distances are the power of $2$. That said, Gray Code ensures the consistency of relative distance representations for every $2^{n}$ ($n\geq 0$) relative distance.
@@
where $G(\cdot)$ represents the function that converts integers into their binary Gray Code representations. The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of $\mathbf{Q}$ and $\mathbf{K}$. $\parallel$ denotes concatenation on the channel dimension $D$.
Notably, the binary nature of Gray Code (comprising only 00 and $1$) aligns intrinsically with the spike-based computation paradigm, avoiding the need for floating-point operations that impose significant implementation overhead on neuromorphic hardware.
@@ 4.4 Log-PE
Here, $\lceil$.$\rceil$ . ⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ is position indices.
Figure 2 (c) shows an illustration of Log-PE. Since the original $\mathbf{AttnMap}$ is a matrix composed of non-negative integers, we aim to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention. Theoretically, if we set the $R_{i,j}$ as $\frac{L-1}{|i-j|+1}$, we could obtain a complete RPE for the spiking Transformers. However, we choose not to pursue this solution, because for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|+1}$ would catastrophically overshadow the original spiking attention activations (See Appendix B). Therefore, using the logarithmic form $R_{i,j}$ represents a compromise that balances the values between the spiking attention map and complete-RPE, while partially capturing relative position information.
@@ 5.1 Datasets
Table 1: Experimental results of time-series forecasting on $4$ benchmarks with various prediction lengths $6,24,48,96$. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. “w/” denotes “with”. The best results for each series of spiking Transformers are highlighted in bold font. $\uparrow$ ($\downarrow$) indicates that the higher (lower) the better. Results highlighted with shading are ours. All results are averaged across $3$ random seeds.
@@
Spike Type $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$
Transformer w/ RoPE R R2$\uparrow$ .$\bf.729$$\bf.560$$\bf.416$$\bf.306$$\bf.787$$730$$\bf.694$$\bf.676$$951$$854$$\bf.763$$\bf.720$$\bf.984$$\bf.978$$\bf.974$$\bf.968$$\bf.756$
@@ 5.2 Time-Series Forecasting
We follow the SeqSNN [6] framework to conduct time-series forecasting experiments. Specifically, we take Spikformer [4], Spike-driven Transformer (SDT) V1 [5], and the current visual state-of-the-art (SOTA) model, QKFormer [20], as the backbone architectures. We modify the SSA mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR. SDT adopts a variant of SSA, which makes it only able to integrate Log-PE but not Gray-PE. We present the performance of the compared SNN models with various positional encoding methods in Table 1. The key findings are as follows:
(1) Directly applying RPE methods to spiking Transformers is ineffective. Specifically, Spikformer with rotary position embedding (RoPE) exhibits poor performance across all benchmarks. As discussed in Section 1, we argue that this limitation stems from the binary nature of spiking neurons during the computation of $\mathbf{Q}$ and $\mathbf{K}$, which makes it difficult to disentangle positional information from sparse spiking activations.
@@ 5.3 Text Classification
Table 2: Accuracy (%) on $6$ text classification benchmarks. Note that QKFormers fail to converge in the text classification task. Experimental results are averaged across $5$ random seeds.
@@
Fine-tuned BERT A $109.8$ $\bm{87.63}$±$\pm 0.18$ $\bm{92.31}$±$\pm 0.17$ $\bm{95.90}$±$\pm 0.16$ $\bm{50.41}$±$\pm 0.13$ $\bm{89.48}$±$\pm 0.16$ $\bm{90.27}$±$\pm 0.13$ $\bm{84.33}$
Spikformer w/o PE $109.8$ $75.87$±$\pm 0.35$ $81.71$±$\pm 0.31$ $91.60$±$\pm 0.30$ $41.84$±$\pm 0.39$ $85.62$±$\pm 0.25$ $86.87$±$\pm 0.28$ $77.25$
Spikformer w/ CPG-PE A $110.4$ $82.42$±$\pm 0.42$ $82.90$±$\pm 0.33$ $92.50$±$\pm 0.25$ $43.62$±$\pm 0.36$ $86.54$±$\pm 0.26$ $\bf 88.49$±$\pm 0.29$ $79.41$
Spikformer-XNOR w/o PE $109.8$ $75.80$±$\pm 0.40$ $81.74$±$\pm 0.40$ $91.50$±$\pm 0.29$ $41.88$±$\pm 0.38$ $85.64$±$\pm 0.31$ $86.66$±$\pm 0.33$ $77.20$
Spikformer-XNOR w/ Gray-PE R $109.8$ $83.73$±$\pm 0.45$ $84.52$±$\pm 0.39$ $92.50$±$\pm 0.33$ $44.06$±$\pm 0.48$ $87.41$±$\pm 0.36$ $88.40$±$\pm 0.30$ $80.11$
Spikformer-XNOR w/ Log-PE R $109.8$ $\bf 83.88$±$\pm 0.40$ $\bf 84.64$±$\pm 0.37$ $\bf 92.80$±$\pm 0.30$ $\bf 44.52$±$\pm 0.43$ $\bf 87.95$±$\pm 0.34$ $88.46$±$\pm 0.28$ $\bf 80.38$
@@ 5.4 Patch-based Image Classification
Table 3: Accuracy (%) on image classification benchmarks. Numbers with denote our implementations. The best and second-best results are highlighted in bold and underlined formats, respectively. The results with shading are ours. Results are averaged across $4$ random seeds.
@@
Furthermore, we present the image classification performance of the state-of-the-art QKFormer integrated with our proposed RPE methods in Appendix C.
@@ 5.5 Capability of Processing Long Sequences
Table 4: Accuracy (%) on $2$ long text classification benchmarks. We set the sentence length to $1024$ for AGNEWS and $2048$ for IMDB.
@@ 5.6 Analysis and Ablation
Consider that: If the number of bits used for encoding relative positions in Gray Code is $b$, then the total number of unique encodings possible is $2^{b}$. We set the maximum sequence length is $L$, so relative distances range from 00 to $L-1$. According to the pigeonhole principle, if $L-1>2^{b}$, there will be at least two distances that are represented identically. This issue can be mitigated by increasing $b$ to cover the range of relative distances up to $L-1$. From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length $=168$), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information. However, for shorter datasets like Metr-la (Length $=12$) and ChnSenti (Length $=32$), $5$ bits are sufficient.
@@ 6 Conclusion
In this work, we have designed several RPE methods for spike Transformers. Our approach preserves the spiking nature of SNNs while effectively representing relative positions. Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements. These empirical results, together with theoretical analysis of the proposed RPE methods, highlight the potential to enhance the versatility and applicability of SNNs across various domains. Future work and limitations are discussed in Appendix E.
@@ Appendix A Proof of Theorem 1
This section provides a detailed mathematical proof of Theorem 1. We use the standard Reflected Binary Code (RBC) $G(x)$ as our Gray Code:
@@
This value has exactly two bits set (at positions $n$ and $n-1$), so the Hamming distance is 2.
@@
Then adding $2^{n}$ to $a$ causes a carry from bit $n$ upwards. Let $c$ be the smallest index greater than $n$ such that bit $c$ in $a$ is 00; bits $n$ through $c-1$ are all $1$. Then:
@@
has ones at bits $n-1$ through $c-1$. Thus,
@@
has ones at only positions $c$ and $n-1$, and all other bits are canceled due to alignment. The result has exactly two bits set, so the Hamming distance is $2$.
@@ Appendix B Ablation Study on Log-PE and XNOR
In this section, we compare the performance of vanilla spiking Transformers with Log-PE, XNOR variants with Log-PE, and models equipped with complete relative positional encoding (C-RPE). As discussed in Section 4.4, C-RPE is implemented by setting $R_{i,j}=\frac{L-1}{|i-j|+1}$ and adding it directly to the attention scores.
@@
Model (Prediction Length $=24$) Metr-la ($L=12$) Electricity ($L=168$)
@@
Spikformer w/ Log-PE .$484$$763$$710^{*}.$ $1.03^{*}$
Spikformer-XNOR w/ Log-PE .$\bf.535$$\bf.719$$\bf.974$$\bf.300$
Spikformer-XNOR w/ C-RPE .$158^{*}.$$967^{*}.$$710^{*}.$ $1.03^{*}$
QKFormer w/ Log-PE .$475$$788$$710^{*}.$ $1.03^{*}$
QKFormer-XNOR w/ Log-PE .$\bf.541$$\bf.715$$\bf.974$$\bf.285$
QKFormer-XNOR w/ C-RPE .$430$$824$$710^{*}.$ $1.03^{*}$
As shown in Table S1, both Spikformer and QKFormer with C-RPE perform significantly worse than their Log-PE counterparts, with some variants failing to converge entirely. This degradation is attributed to the overly large positional encodings disrupting the training dynamics. Furthermore, observed from vanilla Spikformer with Log-PE, we confirm that the dot product, which does not use Hamming distance to measure relative distance, is not suitable for RPE methods.

### v3

@@ Abstract
Spiking neural networks (SNNs) are bio-inspired networks that mimic how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities. SNNs with self-attention mechanisms (spiking Transformers) have recently shown great advancements in various tasks, and inspired by traditional Transformers, several studies have demonstrated that spiking absolute positional encoding can help capture sequential relationships for input data, enhancing the capabilities of spiking Transformers for tasks such as sequential modeling and image classification. However, how to incorporate relative positional information into SNNs remains a challenge. In this paper, we introduce several strategies to approximate relative positional encoding (RPE) in spiking Transformers while preserving the binary nature of spikes. Firstly, we formally prove that encoding relative distances with Gray Code ensures that the binary representations of positional indices maintain a constant Hamming distance whenever their decimal values differ by a power of two, and we propose Gray-PE based on this property. In addition, we propose another RPE method called Log-PE, which combines the logarithmic form of the relative distance matrix directly into the spiking attention map. Furthermore, we extend our RPE methods to a two-dimensional form, making them suitable for processing image patches. We evaluate our RPE methods on various tasks, including time series forecasting, text classification, and patch-based image classification, and the experimental results demonstrate a satisfying performance gain by incorporating our RPE methods across many architectures. Our results provide fresh perspectives on designing spiking Transformers to advance their sequential modeling capability, thereby expanding their applicability across various domains. Our code is available at https://github.com/microsoft/SeqSNN.
@@ 3.2 Spiking Self-Attention
As a result, the attention map $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers. The outputs of the SSA, as well as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, are all spike matrices containing only values of 0 and $1$. The parameters $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}$ are all learnable parameters.
@@ 3.3 Relative Positional Encoding
Here, $\mathbf{R}_{i,j}$ represents the relative positional bias between the $i$-th query and the $j$-th key positions.
Beyond additive bias terms, another widely adopted form of RPE leverages relative positional embeddings directly in the attention computation, where query–position and key–position interactions are parameterized separately. For example, RoPE [14] can be expressed as
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\underbrace{\text{Softmax}\left(\frac{(\mathbf{Q}\mathbf{R}_{i})\cdot(\mathbf{K}\mathbf{R}_{j})^{\operatorname{T}}}{\sqrt{d_{k}}}\right)}_{\mathbf{AttnMap}}\cdot\mathbf{V},$ (6)
where $\mathbf{R}_{i}$ and $\mathbf{R}_{j}$ are position-dependent rotation operators applied to the $i$-th query and $j$-th key vectors, respectively.
A critical aspect of RPE is its adherence to distance consistency: the magnitude of $\mathbf{R}_{i,j}$ is determined exclusively by the relative positional offset |ij||i-j|, ensuring that the model systematically differentiates between proximal and distant tokens. This property enhances the model’s capacity to capture long-range dependencies and generalize across variations in sequence length and structure.
@@ 3.4 Hamming Distance
$d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),\quad\text{where}\quad\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (7)
@@ 3.5 Gray Code
$G(x)=x\oplus\left(x\gg 1\right),$ (8)
@@ 4.3 Gray-PE
We propose that the Gray Code can serve as an approximate approach to relative positional encoding for spiking Transformers. This is supported by the following Theorem˜1:
@@
As illustrated in Figure 2 (b), the Hamming distance $d_{H}(0,1)$ and $d_{H}(1,2)$ both equal $1$ because their relative distance is $1$, i.e., $2^{n},n=0$. Similarly, $d_{H}(0,2)=d_{H}(1,3)$, and $d_{H}(0,4)=d_{H}(1,5)$, as their relative distances are the power of $2$. That said, Gray Code ensures the consistency of relative distance representations for every $2^{n}$ ($n\geq 0$) relative distance.
@@
where $G(\cdot)$ represents the function that converts integers into their binary Gray Code representations. The vector $\bm{l}$ denotes an array of position indexes, specifically $0,1,2,\dots,L-1]$, where $L$ is the sequence length of $\mathbf{Q}$ and $\mathbf{K}$. $\parallel$ denotes concatenation on the channel dimension $D$.
Notably, the binary nature of Gray Code (comprising only 0 and $1$) aligns intrinsically with the spike-based computation paradigm, avoiding the need for floating-point operations that impose significant implementation overhead on neuromorphic hardware.
@@ 4.4 Log-PE
Here, $\lceil.\rceil$ denotes the round-up function, $L$ is the sequence length, and $i,j$ is position indices.
Figure 2 (c) shows an illustration of Log-PE. Since the original $\mathbf{AttnMap}$ is a matrix composed of non-negative integers, we aim to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention. Theoretically, if we set the $R_{i,j}$ as $\frac{L-1}{|i-j|+1}$, we could obtain a complete RPE for the spiking Transformers. However, we choose not to pursue this solution, because for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|+1}$ would catastrophically overshadow the original spiking attention activations (See Appendix˜B). Therefore, using the logarithmic form $R_{i,j}$ represents a compromise that balances the values between the spiking attention map and complete-RPE, while partially capturing relative position information.
@@ 5.1 Datasets
Table 1: Experimental results of time-series forecasting on 44 benchmarks with various prediction lengths $6,24,48,96$. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. “w/” denotes “with”. The best results for each series of spiking Transformers are highlighted in bold font. $\uparrow$ ($\downarrow$) indicates that the higher (lower) the better. Results highlighted with shading are ours. All results are averaged across 33 random seeds.
@@
Spike Type 66 $24$ $48$ $96$ 66 $24$ $48$ $96$ 66 $24$ $48$ $96$ 66 $24$ $48$ $96$
@@ 5.2 Time-Series Forecasting
We follow the SeqSNN [6] framework to conduct time-series forecasting experiments. Specifically, we take Spikformer [4], Spikingformer [31], Spike-driven Transformer (SDT) V1 [5], and the current visual state-of-the-art (SOTA) model, QKFormer [20], as the backbone architectures. We modify the SSA mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR. SDT adopts a variant of SSA, which makes it only able to integrate Log-PE but not Gray-PE. We present the performance of the compared SNN models with various positional encoding methods in Table 1. The key findings are as follows:
(1) Directly applying RPE methods to spiking Transformers is ineffective. Specifically, Spikformers that are directly equipped with RoPE or ALiBi exhibit poor performance across all benchmarks. As discussed in Section 1, we argue that this limitation stems from the binary nature of spiking neurons during the computation of $\mathbf{Q}$ and $\mathbf{K}$, which makes it difficult to disentangle positional information from sparse spiking activations.
@@ 5.3 Text Classification
Table 2: Accuracy (%) on 66 text classification benchmarks. Note that QKFormers fail to converge in the text classification task. Experimental results are averaged across 55 random seeds.
@@
Fine-tuned BERT A $109.8$ $\bm{87.63}$$\pm 0.18$ $\bm{92.31}$$\pm 0.17$ $\bm{95.90}$$\pm 0.16$ $\bm{50.41}$$\pm 0.13$ $\bm{89.48}$$\pm 0.16$ $\bm{90.27}$$\pm 0.13$ $\bm{84.33}$
[…]
@@ 5.4 Patch-based Image Classification
Table 3: Accuracy (%) on image classification benchmarks. Numbers with denote our implementations. The best and second-best results are highlighted in bold and underlined formats, respectively. The results with shading are ours. Results are averaged across 44 random seeds.
@@
Furthermore, we present the image classification performance of the state-of-the-art QKFormer integrated with our proposed RPE methods in Appendix˜C.
@@ 5.5 Capability of Processing Long Sequences
Table 4: Accuracy (%) on 22 long text classification benchmarks. We set the sentence length to $1024$ for AGNEWS and $2048$ for IMDB.
@@
5.6 Discussion on Hardware-Friendliness and Computing Efficiency
Table 5: Evaluation of both time consumption and GPU memory usage for SNNs on Electricity dataset.
Model Time Consumption GPU Memory Usage
s/epoch MB
Spikformer (Original) $137.48$ $10572.56$
Spikformer-XNOR $139.66$ $10608.32$
Spikformer w/ CPG-PE $140.56$ $10963.88$
Although traditional SSA benefits from highly optimized matrix multiplication (GEMM) on GPUs, we would like to clarify that our XNOR-based SSA also retains computational efficiency for the following reasons: First, the core of XNOR-based SSA relies on XNOR and bit-count operations, which are natively supported by digital hardware and neuromorphic processors. These are much cheaper than floating-point multiplications and additions in terms of energy and hardware complexity. Secondly, many neuromorphic accelerators (e.g., Loihi [35], TrueNorth [36]) natively support spike-based bitwise logic, making our XNOR mechanism better aligned with the target deployment platform than conventional floating-point matrix products. Lastly, while matrix multiplication benefits from BLAS acceleration, XNOR and summation over dimensions are also highly parallelizable, and can be efficiently implemented using tensor intrinsics (e.g., bitwise_xnor, popcount, reduce_sum).
We benchmarked both time consumption and GPU memory usage for SNNs in a time-series forecasting task, mainly on the Electricity dataset with 24 of horizon length, as shown in 5. For more analysis on the hardware-friendliness of Log-PE, please refer to the Appendix˜E.
@@ 5.7 Analysis and Ablation
Consider that: If the number of bits used for encoding relative positions in Gray Code is $b$, then the total number of unique encodings possible is $2^{b}$. We set the maximum sequence length is $L$, so relative distances range from 0 to L1L-1. According to the pigeonhole principle, if $L-1>2^{b}$, there will be at least two distances that are represented identically. This issue can be mitigated by increasing $b$ to cover the range of relative distances up to L1L-1. From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length =168=168), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information. However, for shorter datasets like Metr-la (Length =12=12) and ChnSenti (Length =32=32), 55 bits are sufficient.
@@ 6 Conclusion
In this work, we have designed several RPE methods for spike Transformers. Our approach preserves the spiking nature of SNNs while effectively representing relative positions. Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements. These empirical results, together with theoretical analysis of the proposed RPE methods, highlight the potential to enhance the versatility and applicability of SNNs across various domains. Future work and limitations are discussed in Appendix F.
Broader Impact
This work aims to advance the field of spiking neural networks. We hope our work can open new avenues for embedding relative positional encoding within SNNs, thereby expanding their applicability across a wide range of domains. We do not see negative societal impacts of this work.
Acknowledge
The authors would like to thank the anonymous reviewers for their valuable comments. This work was partially supported by the National Natural Science Foundation of China (No. 62076068).
@@ Appendix A Proof of Theorem 1
This section provides a detailed mathematical proof of Theorem˜1. We use the standard Reflected Binary Code (RBC) $G(x)$ as our Gray Code:
@@
This value has exactly two bits set (at positions $n$ and n1n-1), so the Hamming distance is 2.
@@
Then adding $2^{n}$ to $a$ causes a carry from bit $n$ upwards. Let $c$ be the smallest index greater than $n$ such that bit $c$ in $a$ is 0; bits $n$ through c1c-1 are all $1$. Then:
@@
has ones at bits n1n-1 through c1c-1. Thus,
@@
has ones at only positions $c$ and n1n-1, and all other bits are canceled due to alignment. The result has exactly two bits set, so the Hamming distance is $2$.
@@ Appendix B Ablation Study on Log-PE and XNOR
In this section, we compare the performance of vanilla spiking Transformers with Log-PE, XNOR variants with Log-PE, and models equipped with complete relative positional encoding (C-RPE). As discussed in Section˜4.4, C-RPE is implemented by setting $R_{i,j}=\frac{L-1}{|i-j|+1}$ and adding it directly to the attention scores.
@@
Model (Prediction Length =24=24) Metr-la ($L=12$) Electricity ($L=168$)
@@
Spikformer w/ Log-PE .484.484 .763.763 $710^{*}$ $1.03^{*}$
Spikformer-XNOR w/ Log-PE $\bf.535$ $\bf.719$ $\bf.974$ $\bf.300$
Spikformer-XNOR w/ C-RPE $158^{*} .$$967^{*} .$$710^{*}$ $1.03^{*}$
QKFormer w/ Log-PE .475.475 .788.788 $710^{*}$ $1.03^{*}$
QKFormer-XNOR w/ Log-PE $\bf.541$ $\bf.715$ $\bf.974$ $\bf.285$
QKFormer-XNOR w/ C-RPE .430.430 .824.824 $710^{*}$ $1.03^{*}$
As shown in Table˜S1, both Spikformer and QKFormer with C-RPE perform significantly worse than their Log-PE counterparts, with some variants failing to converge entirely. This degradation is attributed to the overly large positional encodings disrupting the training dynamics. Furthermore, observed from vanilla Spikformer with Log-PE, we confirm that the dot product, which does not use Hamming distance to measure relative distance, is not suitable for RPE methods.